ns twistedlog Manual

Edition : 2012.02.16

OpenXM.org

1 Function Manual

1.1 Outline

1.2 Notation

1.3 Functions

1.3.1 ns_twistedlog.twisted_log_cohomology
ns_twistedlog.twisted_log_cohomology(FL,PL,VL)

:: It returns a basis of the middle dimentional twisted logarithmic cohomology group.

FL A list of polynomials.

PL A list of the parameters.

VL A list of the variables.

• Details have not been written. See Japanse documents and examples.

[0] import("ns_twistedlog.rr");
[1] ns_twistedlog.twisted_log_cohomology([x,y,1-x-y],[a,b,c],[x,y]);
-- nd_weyl_gr :0.003848sec(0.008291sec)
-- weyl_minipoly_by_elim :0.006988sec(0.007177sec)
-- generic_bfct_and_gr :0.01325sec(0.02175sec)
generic bfct : [[-1,1],[s,1],[s+a+b+c-1,1]]
S0 : 0
B_S0 length : 1
-- fctr(BF) + base :0.001454sec(0.005543sec)
dimension : 1
[1]
[2] ns_twistedlog.twisted_log_cohomology([x,y,1-x-y],[-1,-2,-3],[x,y]);
-- nd_weyl_gr :0.001845sec(0.001838sec)
-- weyl_minipoly_by_elim :0.003972sec(0.003971sec)
-- generic_bfct_and_gr :0.007363sec(0.007584sec)
generic bfct : [[-1,1],[s,1],[s-7,1]]
S0 : 7
B_S0 length : 36
-- fctr(BF) + base :0.02438sec(0.03323sec)
dimension : 3
[y^2*x^5,y^7,1]
[3] ns_twistedlog.twisted_log_cohomology([x*z+y,x^4+y^5+x*y^4],[0,0],[x,y,z]);
-- nd_weyl_gr :0.004sec(0.0028sec)

weyl_minipoly_by_elim : b-function does not exist
stopped in weyl_minipoly_by_elim2 at line 378 in file "/usr/local/ox/OpenXM/src/asir-
contrib/packages/src/nk_restriction.rr"
378                     error("weyl_minipoly_by_elim : b-function does not exist");
(debug)

Index 3

Index

(Index is nonexistent) (Index is nonexistent)

Short Contents

1 Function Manual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Table of Contents

1 Function Manual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Outline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

   1.3.1 ns_twistedlog.twisted_log_cohomology . . . . . . . . . . . . . . . . 1

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3